# Inferring State Machine from the Protocol Implementation via Large Language Model

### v3

@@ V Experimental Evaluation
To evaluate the effectiveness of  ProtocolGPT, our assessment seeks to answer the following research questions:
@@
RQ1: How effective is ProtocolGPT in inferring state machines compared to other methods?
RQ2: What the impact of the LLM augmentation on the performance of ProtocolGPT?
@@
RQ3: What discrepancies can ProtocolGPT find between implementations of the same protocol?
RQ4: How much more code coverage does fuzzer enhanced by ProtocolGPT achieve compared to baselines?
ProtocolGPT is designed to be compatible with a broad range of LLMs, including offerings from OpenAI [53], Hugging Face [56], etc. For evaluation purposes, we select the widely acclaimed GPT-4 model, renowned for its extensive 1.7 trillion parameters. To provide a consistent and controlled environment for our experiments, they were conducted on a machine boasting 32GB of memory and a 12th-Gen Intel(R) Core(TM) i7-12700 CPU, under Ubuntu 20.04.
@@
Benchmark. Our benchmark encompasses six network protocol implementations that are widely used in the Internet of Things, spanning a broad range of categories. These protocols vary in format and security levels, covering text-based protocols such as RTSP, binary protocols like IKEv2, encrypted protocols like TLS 1.2, and non-encrypted ones like BGP. For each protocol, we identify the implementation with explicitly defined states on GitHub and select the one with the highest number of stars, as shown in Table III. For each protocol in our benchmark, we recruited three experts to independently audit the protocol’s repository code and summarize the state machines. Subsequently, the experts collaborated to discuss and refine the proposed state machines, ultimately consolidating a reliable ground truth. The final outcome serves as the ground truth, with the experts dedicating over 72 hours to this task. To minimize bias, the experts we selected for each protocol were highly familiar with the corresponding protocol RFC. Furthermore, we made it clear to each expert that the ground truth should be based solely on objective analysis, excluding any subjective evaluations. Note that manually constructing ground truth from protocol specifications and source code is a common practice in protocol reverse engineering, as demonstrated by prior works such as RFCNLP[9], NetPlier[13], and StateLifter[10].
@@
Metrics. We focused on the validity of state transitions as it reflects the validity of the states. Therefore, state transitions are categorized into four types: $correct$, $partially$ $correct$, $incorrect$, and $not$ $found$. A state transition $T$ is defined as $\{S_{i},M,S_{t}\}$, where $S_{i}$ is the initial state, $M$ is the message triggering the state transition and $S_{t}$ is the target state. If $S_{i}$, $M$, and $S_{t}$ are all correct, then $T$ is classified as $correct$. If one of $S_{i}$, $M$, or $S_{t}$ is incorrect, then $T$ is $partially$ $correct$. If more than two elements of $T$ are incorrect, then $T$ is classified as $incorrect$. Undetected state transitions are marked as $not$ $found$. We then respectively compute the precision and recall of the inferred FSM as follows:
@@ V-A Effectiveness
TABLE IV: The precision (P) and recall (R) of the state transitions inferred by ProtocolGPT, GPT-4, RFCNLP, Netzob, and NetPlier.
Protocols ProtocolGPT GPT-4 RFCNLP Netzob NetPlier
@@
To evaluate the effectiveness of ProtocolGPT, we select three popular protocol reverse engineering tools and a pre-trained model as baselines: RFCNLP[9], NetZob[57], NetPlier[13], and GPT-4[58]. RFCNLP, NetZob, and NetPlier represent the state-of-the-art in protocol reverse engineering and have been published in leading academic conferences. To compare the performance of augmented model with pre-trained model, we included GPT-4 as an additional baseline. For FSM inference with GPT-4, no source code hints were provided; instead, we directly used the prompt from Section IV-B to generate state machines. As for static analysis technique, We fail to find any open-source static analysis methods capable of inferring state transitions. Existing static analysis techniques are limited to inferring protocol formats.
TABLE V: State machines inferred by ProtocolGPT, GPT-4, RFCNLP, Netzob, and NetPlier. S and T denote the number of states and transitions, respectively; C indicates the number of correct state transitions; PC represents the number of partially correct state transitions; IC refers to the number of incorrect state transitions; and NF signifies the number of state transitions in the ground truth that were not detected by each approach.
@@
We conducted a analysis of the state machines inferred by  ProtocolGPT, GPT-4, RFCNLP, Netzob, NetPlier . Table IV shows the precision and recall of the FSMs inferred by our approach ProtocolGPT versus four base lines GPT-4, RFCNLP, Netzob and NetPlier. All results are the mean of 10 runs and statistically significant with maximum p values less than 0.05. The result demonstrates that the average precision and recall of state transitions inferred by ProtocolGPT achieve 91.42% and 87.09%, while the others often generate over 50% false or miss 70% true transitions.
Analysis of State Transitions. To conduct a comprehensive analysis of the states and state transitions inferred by the five tools, we compile the counts of states and categorize transitions as correct, partially correct, incorrect, and not found, as shown in Table V. GPT-4 generates a substantial number of incorrect transitions and fails to identify many correct ones, whereas ProtocolGPT demonstrates superior performance in this regard. This contrast highlights the potential of LLM augmentation to effectively mitigate hallucinations produced by pre-trained models when inferring FSM. Our observations indicate that RFCNLP which is based on protocol specifications tends to exhibit high precision but low recall. For instance, in the case of the IKEv2 protocol, all four state transitions inferred by RFCNLP are correct. However, the number of transitions was markedly lower than the 23 transitions in the ground truth. This can be attributed to the strong dependence of machine learning-based approaches on the quality of the labeled data. Additionally, the presence of ambiguous definitions in the RFCs can lead to inconsistencies between protocol implementations and the RFCs themselves. Furthermore, the dynamic analysis methods Netzob and NetPlier exhibit a recall of merely 28%. For instance, in the case of the BGP protocol, Netzob successfully identifies 5 out of 88 possible state transitions, whereas NetPlier identifies 6. The inadequate recall of these dynamic methods can be ascribed to their significant dependence on the diversity of input samples. In summary, in comparison to the baseline, our method demonstrates the capacity to accurately infer reliable states and state transitions from the protocol implementations.
@@ V-B Ablation Study
The LLM augmentation process incorporates three primary strategies: code filtering, syntax-aware code segmentation, and vector store. To evaluate the individual contributions of these strategies, we conducted an ablation study. To this end, we develop five distinct variants of ProtocolGPT:
  • V0: GPT-4, with all strategies disabled.
  • V1: V0 with the embedding model.
@@
  • V3: V2 with the code filter.
  • clarify GPT with QA and with original input!!!
@@
Table  VI shows the results in terms of precision and recall. Compared to V0, V3 achieves a 55.59% increase in precision and a 62.95% increase in recall for inferred state machines. This indicates that LLM augmentation plays a crucial role in protocol state machine inference.
Strategy 1: Vector Store. When comparing V1 with V0, there was a noticeable decrease in both precision and recall for V1. We attribute this decline to the different dependent data used by V0 and V1. Specifically, V0 exclusively utilized the pre-trained model, while V1 employed an augmented LLM. During state machine inferring, the output of V0 was derived from the extensive, training data of GPT-4, whereas V1’s augmented LLM relied solely on specific knowledge drawn from external datasets. Notably, V1 do not segment the source code but instead embedded it directly into high-dimensional vectors. The excessive number of tokens within these vectors significantly degraded the model’s performance. For example, during the inference of the L2TP state machine, V1 failed to function properly due to the token count far exceeding the context window of generative model.
Strategy 2: Syntax-aware Code Segmentation. A comparison between V2 and V1 reveals the substantial impact of the syntax-aware code splitting strategy. Specifically, V2 demonstrated an average increase of 27.77% in precision and 12.28% in recall, underscoring the significant benefits of this approach for state machine inference. This improvement is primarily due to the syntax-aware code splitting mechanism, which segments large portions of source code into smaller, more tractable units while preserving the integrity of the syntactic structure. This approach ensures that semantic information is retained, effectively mitigating the risk of meaning loss during the partitioning process. It effectively addresses the problem of token counts exceeding the context window of the generative model. By alleviating this constraint, the overall performance of the model is significantly enhanced.
Strategy 3: Code Filter. The comparison between V3 and V2 highlights the impact of the code filtering mechanism. V3 exhibited a substantial increase in precision and recall, by 41.02% and 55.57%, respectively. This improvement is largely attributable to the code filter’s ability to eliminate code that is irrelevant to the state machine, thus reducing the search space for the retrieval engine. By restricting the augmented LLM from retrieving unrelated vectors, the efficiency of state machine inference is significantly enhanced. However, for protocols with smaller codebases, the effect of the code filter is less pronounced. For example, when inferring the RTSP state machine, the relatively compact size of the protocol’s implementation resulted in minimal differences between the filtered and unfiltered code. As a result, the performance gains of V3 over V2 in this context are marginal.
V-C Inconsistencies Between Protocol Implementations
TABLE VII: The state machines of four IKEv2 protocol implementations inferred by ProtocolGPT
@@ V-C Inconsistencies Between Protocol Implementations
To assess ProtocolGPT’s ability to discern variations between different protocol implementations, we applied it to infer four distinct implementations of the IKEv2 protocol which is a critical component of IPSec suite. A detailed comparative analysis of these implementations is provided in Table VII. Our evaluation revealed substantial discrepancies in the number of states and state transitions among the various IKEv2 implementations, highlighting the nuanced differences in their operational behaviors. For example, strongSwan[41] is characterized by having just 23 state transitions, in contrast to libopenikev2[43], which exhibits as many as 43 state transitions. This notable difference in count of state transition underscores the varying complexity and operational intricacies between these IKEv2 implementations.
Our examination of the state transition relationships in FSMs reveals distinct differences among various protocol implementations. For example, in IKEv2, when the communicating parties complete the establishment of an IKE SA, different protocol implementations exhibit varied state transitions upon the creation of a Child SA. Figure LABEL:fig:state_differenc demonstrate the state transition processes for creating a Child SA in strongSwan [41] and libopenikev2 [43], respectively.
@@
Figure LABEL:fig:strongswan illustrates that in strongSwan, once the peer completes the establishment of the IKE SA (entering the ESTABLISHED state) and initiates a request to create a Child SA with the other peer, the state of the IKE SA remains unchanged. Instead, this action triggers a state transition for the Child SA. In strongSwan, the IKE SA and Child SA manage their state transitions independently. In contrast, as shown in Figure LABEL:fig:libopenikev2, in libopenikev2, after the IKE SA is established (in the STATE_IKE_SA_ESTABLISHED state), sending a request to create a Child SA results in a state transition for the IKE SA to STATE_NEW_CHILD_SA_REQ_SENT. Once the Child SA is created, the state reverts to STATE_IKE_SA_ESTABLISHED.
V-D Security Application
@@
Figure 8: Average line coverage of AFLNet enhanced by  ProtocolGPT, GPT-4, RFCNLP, Netzob, and NetPlier in 5 runs of 24 hours, with a 95% confidence interval.
@@
In protocol fuzzing, state machines are instrumental in exploring more code branches and states effectively. AFLNet leverages protocol response codes to construct protocol state machines, enabling targeted testing of the protocol implementation’s various states as per the state diagram. Utilizing effective message sequences from the corpus as seeds, AFLNet incorporates lightweight mutation algorithms to enhance test coverage. By deriving message sequences from the FSMs inferred by  ProtocolGPT, we generate targeted seeds for AFLNet testing. We adopt the same approach to generate fuzz seeds through the state machines inferred by GPT-4[58], RFCNLP[9], Netzob[57], and NetPlier[13]. To ensure the reliability of our experimental data, each protocol implementation is tested five times, with each session lasting 24 hours, thereby minimizing randomness in the results.
@@
Code Coverage. Figure 8 clearly demonstrates that fuzzers enhanced by  ProtocolGPT outperforms GPT-4, RFCNLP, Netzob and NetPlier by 31.80%, 30.39%, 22.66%, 21.99% in terms of statement coverage, underscoring the capability of  ProtocolGPT-inferred state machines to uncover more code execution paths and states than those inferred by conventional approaches. This is attributed to  ProtocolGPT’s ability to identify a greater number of state transitions, facilitating the discovery of additional code paths. However, in the context of encryption protocols such as IKEv2, TLS1.2, and TLS1.3, AFLNet encounters significant constraints due to its inability to access cryptographic keys during fuzzing, thereby limiting the testing scope to pre-key negotiation phases. The implementation of the BGP protocol represents a distinctive case. Its state transitions, driven by events such as the establishment of connections between autonomous systems, updates to routing information, and administrative operations, differ substantially from those in other protocols. Consequently, message-based seeds alone may prove insufficient to achieve comprehensive code coverage when fuzzing BGP protocol implementations, underscoring the necessity of integrating event-driven triggers into the fuzzing strategy.
@@
Vulnerabilities. The fuzzer enhanced by ProtocolGPT detects two zero-day vulnerabilities and 18 previously known vulnerabilities. In contrast, alternative methods detect only five of the known vulnerabilities. Both zero-day vulnerabilities are present in feng[48], requiring the RTSP protocol server to reach a specific state for successful exploitation. An attacker can exploit these vulnerabilities by sending specially crafted requests, thereby causing the protocol server to crash and executing a Denial-of-Service (DoS) attack. Upon submitting the bug report, the developer promptly confirms the two identified 0-day vulnerabilities. This suggests that the protocol implementation is still active and in use. Consequently, these two 0-day vulnerabilities are likely to present a security threat to industrial software. These detected vulnerabilities can be attributed to our methodology, which facilitates the inference of a reliable and comprehensive protocol state machine, enabling the protocol fuzzer to trigger a wider range of protocol states.
VI Related work
Natural Language Processing Pacheco et al. advanced a novel methodology for inferring protocol state machines, anchored in NLP and adopting a data-driven paradigm [9]. This method involves the training of deep learning models on a comprehensive corpus of natural language data, encompassing technical documentation and protocol specification documents. Subsequently, leveraging labeled data, this approach facilitates the derivation of relationships between variables within the protocol specifications via the model, culminating in the articulation of a finite state machine representative of the protocol.
Dynamic Analysis Kleber et al. unveiled a novel strategy for the categorization of message types within binary protocols [14]. This strategy relies on a similarity metric using continuous value ranges to compare feature vectors based on message fields instead of discrete byte values, thereby facilitating an enhanced discernment of message format patterns. Further contributing to the field, Ye et al. advanced a probabilistic approach to the reversal of network protocols [13]. By constructing a joint distribution between these random variables and the observed values within messages, probabilistic reasoning is employed to ascertain potential keywords. Gopinath et al. introduced a comprehensive algorithm designed to deduce the syntax of inputs from programs, distinguishing itself by eschewing conventional data flow analysis methodologies [15]. Uniquely, this methodology obviates the necessity for heuristic strategies traditionally employed to discern parsing patterns, rendering it universally applicable across all recursive parsers that operate on the basis of program stacks.
Static Analysis Shi et al. pioneered the development of a protocol state machine inference methodology predicated on static analysis [10]. This innovative approach employs static analysis to examine protocol formats derived from parsers within the protocol implementations. The extraction of state machines is achieved by interpreting each iteration of a loop as a distinct state, with the interdependencies among loop iterations being regarded as transitions between states.
VII Conclusion
Inferring state machines from protocol implementations is inherently laden with complexities and challenges. In the context of intricate protocol implementations, conventional methodologies exhibit limitations in accurately inferring state machine. On one hand, static analysis contends with the dilemma of state explosion during the inference process. On the other hand, with respect to the specific endeavor of state machine inference, the outcomes yielded by static analysis lack the requisite precision and comprehensiveness. Consequently, We propose an innovative state machine inference approach powered by augmented LLM and conduct experiments on the benchmark. The outcomes of these experiments underscore the capability of LLMs to facilitate the derivation of relatively robust state machines from protocol implementations. Moreover, employing  ProtocolGPT enabled us to conduct a comparative analysis of state machines across various implementations of identical protocols, through which we successfully delineated discrepancies among the state machines corresponding to these implementations. Conclusively, our investigations corroborated that the state machines inferred by  ProtocolGPT could significantly augment protocol fuzzers by enhancing code coverage and facilitating the uncovering of previously unidentified bugs.

### v4

@@ V Experimental Evaluation
To evaluate the effectiveness of  ProtocolGPT, we try to answer the following research questions:
@@
RQ1: How effective is ProtocolGPT in inferring state machines compared to other methods? (Section A)
RQ2: How the components effect the performance of ProtocolGPT? (Section B)
@@
RQ3: What discrepancies can be found between different implementations of the same protocol? (Section C)
RQ4: How can ProtocolGPT assist find more security bugs? (Section D)
We select the widely acclaimed GPT-4 model (the latest model at the time of our experiment), renowned for its extensive 1.7 trillion parameters. Following the recommendation to use a low temperature for generating factual responses, we set it to 0.2. Our experiments were conducted on a machine boasting 32GB of memory and a 12th-Gen Intel(R) Core(TM) i7-12700 CPU, under Ubuntu 20.04.
@@
Benchmark. Our benchmark contains 6 network protocol implementations that are widely used in the Internet of Things, spanning a broad range of categories. These protocols vary in format and security levels, covering text-based protocols such as RTSP, binary protocols like IKEv2, encrypted protocols like TLS 1.2, and non-encrypted ones like BGP. For each protocol, we identify the implementation with explicitly defined states on GitHub and select the one with the highest number of stars, as shown in Table II. To set up the ground truth, we recruited three knowledgeable experts with over 72 hours human effort, to independently audit the protocols’ repository code and summarize the state machines. Subsequently, the experts collaborated to discuss and refine the proposed state machines, considering a reliable ground truth. To minimize bias, the experts we selected for each protocol were highly familiar with the corresponding protocol RFC. Furthermore, we made it clear to each expert that the ground truth should be based solely on objective analysis, excluding any subjective evaluations. Note that manually constructing ground truth from protocol specifications and source code is a common practice in protocol reverse engineering, as demonstrated by prior works such as RFCNLP[6], NetPlier[9], and StateLifter[7].
@@
Metrics. We focus on the validity of state transitions as it also reflects the validity of the states. State transitions are categorized into four types: $correct$, $partially$ $correct$, $incorrect$, and $not$ $found$. A state transition $T$ is defined as $\{S_{i},M,S_{t}\}$, where $S_{i}$ is the source state, $M$ is the message triggering the transition and $S_{t}$ is the destination state. If $S_{i}$, $M$, and $S_{t}$ are all correct, then $T$ is classified as $correct$. If one of $S_{i}$, $M$, or $S_{t}$ is incorrect, then $T$ is $partially$ $correct$. If more than two elements of $T$ are incorrect, then $T$ is classified as $incorrect$. Undetected state transitions are marked as $not$ $found$. We then respectively compute the precision and recall of the inferred FSM as follows:
@@ V-A Effectiveness
TABLE IV: Precision and recall of the state transitions inferred by ProtocolGPT, GPT-QA, RFCNLP, Netzob, and NetPlier.
Protocols ProtocolGPT GPT-QA RFCNLP Netzob NetPlier
@@
To evaluate the effectiveness of ProtocolGPT, we select three popular protocol reverse engineering tools as baselines: RFCNLP[6], NetZob[35], NetPlier[9]. RFCNLP, NetZob, and NetPlier represent the state-of-the-art in protocol reverse engineering and have been published in leading academic conferences. To compare with pre-trained model, we include GPT-QA as an additional baseline. For FSM inference with GPT-QA, no source code hints were provided. Instead, we directly use the prompt in Section IV-B to generate FSMs using GPT-4. Regarding static analysis technique, We fail to find any open-source tools capable of inferring FSMs. Existing static analysis techniques are limited to inferring protocol formats.
We conduct a analysis of the FSMs inferred by  ProtocolGPT, GPT-QA, RFCNLP, Netzob, NetPlier . Table IV shows the precision and recall of the state transitions. All results are the average of 10 runs and statistically significant with maximum p values less than 0.05. We can observe that the average precision and recall of state transitions inferred by ProtocolGPT achieve 91.42% and 87.09%, outperforming baselines by more than 30% and 55%. Our method significantly outperforms GPT-QA, demonstrating its ability to effectively enhance pre-trained models. Our analysis shows that the learning-based approach, RFCNLP, tends to achieve high precision but low recall, as demonstrated by their more than 10% higher precision than ours in the case of the RTSP protocol. This can be attributed to the strong reliance on the quality of the labeled data. Additionally, the dynamic analysis methods Netzob and NetPlier exhibit a recall of merely 28%, which is 60% lower than that of ProtocolGPT. The limited recall is primarily due to their heavy dependence on the diversity of inputs. These results highlight that our method not only enhances pre-trained models but also overcomes the limitations of conventional approaches.
TABLE V: State machines inferring results. S and T denote the number of states and transitions, respectively. The C, PC, I, and NF represent the correct, partially correct, incorrect, and not found state transitions, respectively.
@@
Analysis of State Transitions. To conduct a comprehensive analysis of the states and state transitions inferred by the five tools, we statistical the details of the result, as shown in Table V. GPT-QA generates a substantial incorrect transitions and fails to identify many correct ones, whereas ProtocolGPT demonstrates superior performance in this regard. This contrast highlights the potential of LLM augmentation to effectively mitigate hallucinations produced by pre-trained models. Although our approach has a 10% lower precision compared to the RFCNLP, it is more robust.Specifically, in the case of RTSP, RFCNLP identify 11 correct transitions, while our method identify 12. The superior performance of our method in other protocols can be attributed to ambiguous definitions in the RFCs. Furthermore, the dynamic analysis methods Netzob and NetPlier often miss substantial correct transitions. For instance, in the case of the BGP protocol, Netzob successfully identifies 5 out of 88 transitions, and NetPlier identifies 6. In summary, compared to the baselines, our method demonstrates the ability to accurately infer reliable FSMs from the protocol implementations.
@@ V-B Ablation Study
Our LLM augmentation method incorporates three primary strategies: code filtering, syntax-aware code segmentation, and code embedding. To evaluate the individual contributions of the strategies, we conducted the following ablation study. Therefore, We develop five variants of ProtocolGPT:
  • GPT-QA: GPT-4 with only the prompt in Section IV-B.
  • V0: Input with original source code via GPT-4.
  • V1: V0 with code embedding.
@@
  • V3: V2 with code filter.
@@
The results are presented in Table  VI. V0 is unable to work on any protocols due to the limitation of model’s context window. The performance of V1, V2, and V3 shows a consistent improvement, indicating that all three components make positive contributions. In comparison to GPT-QA, V3 demonstrates a significant improvement of 55.59% in precision and 62.95% in recall. These results underscore the critical role that LLM augmentation plays in the FSM inference.
Strategy 1: Vector Store. Compared to V0, the improvement in V1 is attributed to the ability to retrieve state information from the embedded code. In contrast, V0 is unable to process the entire codebase. However, when comparing V1 with GPT-QA, there was a noticeable decrease in both precision and recall for V1. We attribute this decline to the different dependent knowledge used by V0 and V1. Specifically, the output of V0 was derived from the training data of GPT-4, whereas V1’s augmented LLM relied on specific knowledge drawn from external datasets.
Strategy 2: Syntax-aware Code Segmentation. A comparison between V2 and V1 reveals the impact of the syntax-aware code segmentation. Specifically, V2 demonstrated an average increase of 27.77% in precision and 12.28% in recall, underscoring the significant benefits of this strategy. This improvement is mainly due to the mechanism splitting the code into more manageable units while preserving the integrity of the syntactical structure. This approach ensures that semantic information is retained, effectively mitigating the risk of meaning loss during the partitioning process. It effectively addresses the limitation of the LLM’s context window.
Strategy 3: Code Filter. The comparison between V3 and V2 highlights the impact of the code filter. V3 exhibited a substantial increase in precision and recall by 41.02% and 55.57%. This improvement is attributable to the code filter’s ability to eliminate code that is irrelevant to FSM, thus reducing the search space for the retrieval. By restricting the augmented LLM from retrieving unrelated data, the efficiency of FSM inference is significantly enhanced. However, the effect of the code filter is less pronounced for concise repository. For example, due to RTSP’s compact implementation, the minimal discrepancies between the filtered and unfiltered code led to marginal performance gains of V3 over V2.
V-C Discrepancies between Protocol Implementations
TABLE VII: The state machines of four IKEv2 protocol implementations inferred by ProtocolGPT.
@@ V-C Discrepancies between Protocol Implementations
To assess ProtocolGPT’s ability to discern variations between different implementations, we applied it to four distinct implementations of the IKEv2 protocol which is a critical component of IPSec suite. A detailed comparative analysis of these implementations is provided in Table VII. Our evaluation revealed discrepancies in the FSMs among the various IKEv2 implementations, highlighting the nuanced discrepancies in their operational behaviors. For example, in IKEv2, when the communicating parties complete the establishment of an IKE SA, different implementations exhibit varied state transitions upon the creation of a Child SA. Figure LABEL:fig:state_differenc demonstrate the state transition processes for creating a Child SA in strongSwan [23] and libopenikev2 [24], respectively.
@@
Figure LABEL:fig:strongswan illustrates that in strongSwan, once the peer completes the establishment of the IKE SA and initiates a request to create a Child SA with the other peer, the state of the IKE SA remains unchanged. Instead, this action triggers a transition for the Child SA. In strongSwan, the IKE SA and Child SA manage their state independently. In contrast, as shown in Figure LABEL:fig:libopenikev2, in libopenikev2, after the IKE SA is established, sending a request to create a Child SA results in a transition for the IKE SA to STATE_NEW_CHILD_SA_REQ_SENT. Once the Child SA is created, the state reverts to STATE_IKE_SA_ESTABLISHED.
V-D Security Applications
@@
Figure 7: Average line coverage of AFLNet enhanced by different methods in 5 runs of 24 hours, with a 95% confidence interval.
@@
In protocol fuzzing, state machines are instrumental in exploring more code branches and states. AFLNet leverages response codes to construct state machines, enabling targeted testing of the protocol’s various states. Utilizing effective message sequences from the corpus as seeds, AFLNet incorporates lightweight mutation algorithms to enhance coverage. By deriving message sequences from the FSMs inferred by  ProtocolGPT, we generate targeted seeds for AFLNet testing. We adopt the same approach to generate fuzz seeds through the FSMs inferred by GPT-QA[36], RFCNLP[6], Netzob[35], and NetPlier[9]. To reduce randomness in the experiment results, we employed fuzzers across all protocols, running five repetitions over 24 hours.
@@
Code Coverage. Figure 7 demonstrates that fuzzers enhanced by  ProtocolGPT outperforms GPT-QA, RFCNLP, Netzob and NetPlier by 31.80%, 30.39%, 22.66%, 21.99% in terms of statement coverage. This is attributed to  ProtocolGPT’s ability to identify a greater number of state transitions, facilitating the discovery of additional code paths. However, in the context of encryption protocols such as IKEv2, TLS1.2, and TLS1.3, AFLNet encounters significant constraints due to its inability to access cryptographic keys during fuzzing, thereby limiting the testing scope to pre-key negotiation phases. The BGP protocol has unique state transitions triggered by events like routing updates. As a result, message-based seeds can not fully cover the code during fuzzing, which highlights the need for event-driven triggers in protocol fuzzing.
@@
Vulnerabilities. The fuzzer enhanced by ProtocolGPT detects two 0-day vulnerabilities and 18 previously known vulnerabilities. In contrast, other methods detect only five of the known vulnerabilities. Both 0-day vulnerabilities are present in Feng’s repository[30], requiring the RTSP server to reach a specific state for successful exploitation. The attackers can exploit these vulnerabilities by sending specially crafted requests, causing the server to crash and executing a Denial-of-Service (DoS) attack. Upon submitting the bug report, the developer promptly confirms the two 0-day vulnerabilities. This suggests that the implementation is still active and in use. Consequently, these two 0-day vulnerabilities are likely to present a security threat to industrial software. These detected vulnerabilities can be attributed to our methodology, which facilitates the inference of a reliable protocol FSM, enabling the fuzzer to trigger a wider range of states.
VI Conclusion
In this paper, we introduce a novel state machine inference method utilizing augmented LLMs and perform benchmark experiments. The results demonstrate the potential of LLMs in deriving robust state machines from protocol implementations. Additionally, the use of  ProtocolGPT allowed for a comparative analysis of state machines across different implementations of the same protocols, revealing discrepancies among them. Ultimately, our findings confirm that the state machines inferred by  ProtocolGPT can substantially enhance protocol fuzzers by improving code coverage and aiding in the discovery of previously undetected bugs.